Packaging Solutions Corporation manufactures and sells a wide variety of packaging products. Performance reports are prepared monthly for each department. The planning budget and flexible budget for the Production Department are based on the following formulas, where q is the number of labor-hours worked in a month:

Cost Formulas
Direct labor $16.60q
Indirect labor $4,400 + $1.50q
Utilities $5,700 + $0.70q
Supplies $1,400 + $0.40q
Equipment depreciation $18,700 + $2.40q
Factory rent $8,100
Property taxes $2,600
Factory administration $13,700 + $0.60q

The Production Department planned to work 4,300 labor-hours in March; however, it actually worked 4,100 labor-hours during the month. Its actual costs incurred in March are listed below:

Actual Cost Incurred in March
Direct labor $ 69,700
Indirect labor $ 10,030
Utilities $ 9,100
Supplies $ 3,330
Equipment depreciation $ 28,540
Factory rent $ 8,500
Property taxes $ 2,600
Factory administration $ 15,510

Required:

1. Prepare the Production Department’s planning budget for the month.

2. Prepare the Production Department’s flexible budget for the month.

3. Prepare the Production Department’s flexible budget performance report for March, including both the spending and activity variances.
Answer #1
1
Planning Budget
Budgeted labor hours 4300
Direct labor 71380 =4300*16.6
Indirect labor 10850 =4400+(4300*1.5)
Utilities 8710 =5700+(4300*0.7)
Supplies 3120 =1400+(4300*0.4)
Equipment Depreciation 29020 =18700+(4300*2.4)
Factory rent 8100
Property taxes 2600
Factory administration 16280 =13700+(4300*0.6)
Total expenses 150060
2
Flexible Budget
Actual labor hours 4100
Direct labor 68060 =4100*16.6
Indirect labor 10550 =4400+(4100*1.5)
Utilities 8570 =5700+(4100*0.7)
Supplies 3040 =1400+(4100*0.4)
Equipment Depreciation 28540 =18700+(4100*2.4)
Factory rent 8100
Property taxes 2600
Factory administration 16160 =13700+(4100*0.6)
Total expenses 145620
3
Actual Spending variance Flexible budget Activity variance Planning budget
Labor hours 4100 4100 4300
Direct labor 69700 1640 U 68060 3320 F 71380
Indirect labor 10030 520 F 10550 300 F 10850
Utilities 9100 530 U 8570 140 F 8710
Supplies 3330 290 U 3040 80 F 3120
Equipment Depreciation 28540 0 None 28540 480 F 29020
Factory rent 8500 400 U 8100 0 None 8100
Property taxes 2600 0 None 2600 0 None 2600
Factory administration 15510 650 F 16160 120 F 16280
Total expenses 147310 1690 U 145620 4440 F 150060